An egg rolls off the edge of a table 1.2-meter-high and strikes the floor at a

point 2 meter horizontally from the edge of the table.


1
Expert's answer
2021-03-02T18:06:35-0500

Answer

Using

For time "t=\sqrt{\frac{2h}{g}}\\=\sqrt{\frac{2\times1.2}{10}}\\=0.49sec"

Now speed of rolls is given by

"v=\frac{distance}{time}=\frac{2}{0.49}=4.08m/s"